OpenStudy (anonymous):

If T : R2 → R2 is a linear transformation such that T1 0 = 2 5 and T0 1 = −1 6 , find T 5 −3. (

OpenStudy (anonymous):

oh wow that didnt work lol It says T(\[\left(\begin{matrix}1 \\ 0\end{matrix}\right) =\left(\begin{matrix}2 \\ 5\end{matrix}\right)\] \[T\left(\begin{matrix}0 \\ 1\end{matrix}\right)=\left(\begin{matrix}-1 \\ 6\end{matrix}\right)\] Find\[T\left(\begin{matrix}5 \\ -3\end{matrix}\right)\]

OpenStudy (jamesj):

Ok, so the first equation gives you first column of matrix; the second equation the second. \[T = \left[\begin{matrix} 2 & -1 \\ 5 & 6 \end{matrix}\right]\] You should convince yourself of that by multiplying back out the two equations. Now that you have T, it should be easy to calculate \[\ T \left(\begin{matrix} 5 \\ -3 \end{matrix}\right)\]

OpenStudy (zarkon):

\[T \left(\begin{matrix} 5 \\ -3 \end{matrix}\right)=5\cdot T \left(\begin{matrix} 1 \\ 0 \end{matrix}\right)+-3\cdot T \left(\begin{matrix} 0 \\ 1 \end{matrix}\right)\]

OpenStudy (jamesj):

Alternatively note that \[\left(\begin{matrix} 5 \\ -3 \end{matrix}\right) = 5 \left(\begin{matrix} 1 \\ 0 \end{matrix}\right) + -3 \left(\begin{matrix}0 \\ -1 \end{matrix}\right)\] So .. and Zarkon just beat me too it!

OpenStudy (anonymous):

so... the answer would be \[\left(\begin{matrix}13\\ 7\end{matrix}\right)\] Its that easy? lol

OpenStudy (anonymous):

If it is that easy my professor needs to stop making it sound so complicated during class

OpenStudy (zarkon):

it is that easy

OpenStudy (anonymous):

Can you help me with some more "easy" ones that arent related to this kind of problem but i need some clarity?